OpenStudy (anonymous):

Write the standard form of the line that is parallel to 2x + 3y = 4 and passes through the point (1, -4).

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

parallel lines should have same slope

OpenStudy (anonymous):

The slope is -2/3

Nnesha (nnesha):

so equation is 2x + 3y =b replace x and y by this order pair ( 1 ,-4) and solve for b

OpenStudy (ybarrap):

Also, here is a visual solution to your problem - https://tube.geogebra.org/student/m837443

OpenStudy (anonymous):

b=-10

Nnesha (nnesha):

yes so 2x + 3y =b replace b by -10 and done!

OpenStudy (anonymous):

So the answer is 2x+3y=-10

Nnesha (nnesha):

yep

OpenStudy (anonymous):

Thanks so much again! I really appreciate it:)

Nnesha (nnesha):

my pleasure :)